OpenStudy (anonymous):

A model for a company’s revenue is r=-15p^2+300p+12,000 ,where p is the price in dollars of the company’s product. What price will maximize revenue?

OpenStudy (anonymous):

\[r=-15p^2+300p+12,000\]

OpenStudy (anonymous):

magic formula is \[-\frac{b}{2a}\]

OpenStudy (anonymous):

in this case a = -15, b = 300 so \[-\frac{b}{2a}=\frac{300}{30}=10\]

OpenStudy (anonymous):

replace p by 10 and you win. \[-15(10)^2+300\times 10+1200=2700\]